# AdaSCALE: Adaptive Scaling for OOD Detection

### v2

@@ 4.2.1 Perturbation Mechanism
Let $x\in\mathbb{R}^{C_{\text{in}}\times H\times W}$ be an input image with $C_{\text{in}}$ input channels, $H$ height, and $W$ width. We denote channel value at position $(c,h,w)$ as $x[c,h,w]$. To identify channel values for perturbation, we employ pixel attribution that quantifies each input element’s influence on the model’s prediction. An attribution function, $AT(x,c,h,w)$, assigns a score to each channel value, with lower absolute scores indicating less influence. We select $o\%$ of channel value indices with lowest absolute attribution scores, forming the set $R$. We use a gradient-based attribution $AT(x,c,h,w)=\frac{\partial(g_{W}(f_{\theta}(x)))_{y_{pred}}}{\partial x[c,h,w]}$, where $y_{pred}$ is predicted class index. To create a perturbed input, we select a subset $R$ containing $o\%$ of channel values to perturb. If $\varepsilon$ is perturbation magnitude, the perturbed image $x^{\varepsilon}$ is obtained as:
$x^{\varepsilon}[c,h,w]=\begin{cases}x[c,h,w]+\varepsilon\cdot\text{sign}(AT(x,c,h,w)),&\text{if }(c,h,w)\in R\\ x[c,h,w],&\text{if }(c,h,w)\notin R\end{cases}$ (2)
Remark: Is gradient attribution necessary for perturbation? While we employ gradient-based attribution for principled pixel selection for perturbation, as we show later in Section˜E.5, it is important to note that even random selection empirically performs similarly, whereas selecting salient pixels degrades performance.
@@ 4.2.2 Activation Shift as OOD Indicator
Figure 3: $Q_{\text{OOD}}$$Q_{\text{ID}}$ vs $Q^{\prime}_{\text{OOD}}$$Q^{\prime}_{\text{ID}}$ in various OOD datasets with ResNet-50 on ImageNet-1k. $Q^{\prime}_{\text{OOD}}/Q^{\prime}_{\text{ID}}>Q_{\text{OOD}}/Q_{\text{ID}}$ suggests $C_{o}$ helps mitigate overconfident estimations.
Indeed, Figure˜3 illustrates that $Q^{\prime}_{\text{OOD}}/Q^{\prime}_{\text{ID}}>Q_{\text{OOD}}/Q_{\text{ID}}$, suggesting that the correction term $C_{o}$ helps mitigate overconfident estimations. If $\bar{Q}_{s}=\{\bar{Q}^{\prime}_{1},\bar{Q}^{\prime}_{2},...,\bar{Q}^{\prime}_{n_{\text{val}}}\}$ be the set of $Q^{\prime}$ values on $n_{\text{val}}$ ID validation samples, we could transform any $Q^{\prime}$ into a normalized probability scale by constructing empirical cumulative distribution function (eCDF) derived from $\bar{Q}_{s}$. The eCDF, denoted as $F_{Q^{\prime}}(Q^{\prime})$, can be defined as: $F_{Q^{\prime}}(Q^{\prime})=\frac{1}{n_{\text{val}}}\sum_{i=1}^{n_{\text{val}}}\mathbbm{1}(\bar{Q}^{\prime}_{i}\leq Q^{\prime})$, where $\mathbbm{1}(\cdot)$ is the indicator function. A higher value of $F_{Q^{\prime}}(Q^{\prime})$ indicates a higher likelihood of the sample being OOD. Importantly, our experiments suggest that as few as 10 ID validation samples are sufficient to construct an effective eCDF for this purpose (See Table˜9).
@@ 4.3 Proposed Approach: Adaptive Scaling
Building on our observations, we propose AdaSCALE (Adaptive SCALE), a novel approach that introduces dynamic, sample-specific adjustments to the scaling procedure. The key insight is that $p^{th}$ percentile threshold should be a function of each test sample’s estimated OODness rather than a fixed value. The scaling factor $r$ increases as the $p^{th}$ percentile threshold rises (i.e., when more activations are excluded from the denominator in Equation 1). For optimal ID-OOD separation, we must scale ID samples more strongly than OOD samples, requiring a higher $p^{th}$ percentile for ID samples. We define an adaptive percentile threshold as: $p=p_{\min}+(1-F_{Q^{\prime}}(Q^{\prime})\cdot(p_{\max}-p_{\min})$, where $p_{\text{min}}$ and $p_{\text{max}}$ are hyperparameters that define the minimum and maximum limits of percentile threshold. It ensures samples with lower estimated OODness receive higher percentile thresholds, resulting in stronger scaling. (See Algorithm˜1). We implement two variants: AdaSCALE-A scales activations as $\mathbf{a}_{\text{scaled}}=\mathbf{a}\cdot\exp(r)$ (Djurisic et al., 2023; Xu et al., 2024). AdaSCALE-L scales logits as $\mathbf{z}_{\text{scaled}}=\mathbf{z}\cdot r^{2}$ (Djurisic et al., 2024). This approach is also outlined in Figure˜4.
@@ 5 Experiments
We use pre-trained models provided by PyTorch for ImageNet-1k experiments. For CIFAR experiments, we train three models per network using the standard cross-entropy loss and report the mean results across these three independent trials. The evaluation setup is provided in Table˜1. Although we tune $(p_{\text{min}},p_{\text{max}})$ for optimal results in each case (see Appendix˜G), near-optimal performance can be achieved by tuning only $p_{\text{max}}$ while fixing $(\lambda,k_{1},k_{2},o,\epsilon,p_{\text{min}})$ to $(10,1\%,5\%,5\%,0.5,60)$ across all 22 cases (See Table˜5). These values were determined via automatic parameter search (Yang et al., 2022; Zhang et al., 2023b) with respect to the AUROC metric using ResNet-50 model on ImageNet-1k benchmark. Best results are bold, and second-best results are underlined.
@@
Baselines. MSP (Hendrycks & Gimpel, 2017), EBO (Liu et al., 2020), ReAct (Sun et al., 2021), MLS (Hendrycks et al., 2022a), ASH (Djurisic et al., 2023), SCALE (Xu et al., 2024), BFAct (Kong & Li, 2022), LTS (Djurisic et al., 2024), OptFS (Zhao et al., 2024). Currently, SCALE is the best-performing method (with ResNet-50), while OptFS is the best-generalizing method.
@@ 5.1 Empirical Results
[…]
CIFAR benchmark. We compare AdaSCALE with post-hoc baselines on CIFAR benchmarks using WRN-28-10 and DenseNet-101 networks, reporting the averaged performance in Table˜3. AdaSCALE outperforms all methods in average AUROC metric across CIFAR benchmarks in both near- and far-OOD detection. For far-OOD detection on CIFAR-10 benchmark, AdaSCALE-A achieves the best FPR@95 score of 33.11, outperforming the MSP baseline by approximately 1.4 points. Similarly, AdaSCALE-A attains the best FPR@95 / AUROC of 43.07 / 90.31 in near-OOD detection, though MSP remains competitive. In near-OOD detection on CIFAR-100 benchmark, AdaSCALE-A achieves the highest AUROC of 81.35, while in far-OOD detection, AdaSCALE-L reaches the best performance with FPR@95 / AUROC of 52.49 / 82.21. While activation-shaping methods perform well in ImageNet-1k, they seem to underperform in CIFAR. In contrast, AdaSCALE achieves consistently superior performance across all setups.
ImageNet-1k benchmark. We compare our proposed method, AdaSCALE, with recent state-of-the-art approaches across eight architectures on the ImageNet-1k benchmark, as presented in Table˜2. AdaSCALE demonstrates consistently strong performance across all architectures compared to existing methods. Specifically, it surpasses the best-generalizing method, OptFS, by $\textbf{14.94\%}/\textbf{8.96\%}$ in the FPR@95/AUROC metric for near-OOD detection across all architectures. Additionally, it outperforms the best-performing method, SCALE (on ResNet-50), by $\textbf{12.96\%}/\textbf{6.44\%}$ in the same metric. A closer observation reveals that while OptFS excels in architectures such as EfficientNet, ViT-B-16, and Swin-B, scaling baselines perform comparably or even better in architectures like ResNet-50, ResNet-101, RegNet-Y-16, and DenseNet-201. In contrast, AdaSCALE-A achieves the best performance in near-OOD detection across all architectures, except for Swin-B, where BFAct performs optimally. Furthermore, effectiveness of AdaSCALE extends beyond near-OOD detection to far-OOD detection, demonstrating an average gain of 21.67% over OptFS in the FPR@95 metric.
Table 4: FSOOD detection results on ImageNet-1k averaged over 8 architectures.
Method Near-OOD Far-OOD FPR@95 $\downarrow$ AUROC $\uparrow$ FPR@95 $\downarrow$ AUROC $\uparrow$ ReAct 87.22 51.38 67.23 69.53 ASH 87.01 52.02 72.36 65.70 SCALE 86.75 52.27 69.36 68.97 BFAct 87.12 51.14 66.13 69.69 LTS 86.46 53.29 66.76 71.63 OptFS 85.83 52.17 63.32 71.44 AdaSCALE-A 81.34 55.03 58.87 72.41 AdaSCALE-L 81.62 55.14 59.19 72.85
FSOOD Detection. FSOOD detection extends conventional OOD detection by incorporating model’s ability to generalize on covariate-shifted ID inputs. We present FSOOD detection results in Table˜4. We can observe that this is a highly challenging task, as covariate-shifted ID datasets cause a significant performance drop for all methods compared to conventional case. Despite this, AdaSCALE outperforms OptFS by 4.49 and 4.13 points on average in the FPR@95 metric for FSOOD detection across both near- and far-OOD datasets.
Accuracy: Like SCALE and LTS, AdaSCALE applies linear transformations to scale activations or logits, preserving accuracy, unlike post-hoc rectification methods (ASH, ReAct).
@@ 5.2 Ablation / hyperparameter studies / Discussion
Sensitivity study. We study the sensitivity of AdaSCALE-A hyperparameters with ResNet-50 on ImageNet-1k (Table˜7). The optimal perturbation magnitude is $\varepsilon=0.5$, close to RGB means and sufficient to disturb peak activations. Setting $\lambda=10$ ensures activation shift $Q$ dominates the regularizer $C_{o}$. Optimal performance at $k_{1}=1\%$ confirms that shifts in peak activations are the most reliable OOD signal, while $k_{2}=5\%$ empirically performs best. $o=5\%$ yielding optimal result supports our hypothesis that perturbation should be trivial enough to only disturb peak activations. (See Appendix˜E for complete sensitivity study.)
@@
Constrained hyperparameter tuning. The hyperparameters $(\lambda,k_{1},k_{2},o,\epsilon)$ = $(10,1\%,5\%,5\%,0.5)$ determined with ResNet-50 model in ImageNet-1k benchmark generalize across architectures and datasets. To demonstrate this, we conduct experiments only allowing $p_{\text{max}}$ to be tuned across 8 and 2 diverse architectures in ImageNet-1k and CIFAR-100 datasets, respectively and present the results in Table˜5. Inspired by early works that tune the percentile hyperparameter in $[60,99]$, we set $p_{\text{min}}$ to the lower limit 60. Even under this highly constrained setting, AdaSCALE significantly outperforms the previous state-of-the-art (OptFS) by an average of 13% (FPR95) / 8% (AUROC) on near-OOD and 14% (FPR95) / 2% (AUROC) on far-OOD datasets on ImageNet-1k benchmark. The superiority of AdaSCALE over OptFS is also evident in a small-scale setting (CIFAR-100).
Estimated OODness $Q^{\prime}$. Adaptive scaling depends on estimated OODness to determine the extent of scaling. We study the effect of various estimated OODness functions using ResNet-50 network (ImageNet-1k) in Table˜6. It clearly shows Q component of $Q^{\prime}$ being most critical while $\sum_{k=1}^{k_{2}}\mathbf{a}^{\varepsilon}_{\text{argsort}(\mathbf{a})_{k}}$ as correction term being a relatively superior choice. However, estimated OODness alone – without adaptive scaling – does not result in strong performance.
@@
Table 8: Robustness to Gaussian blur, JPEG compression, and adversarial training.
Method Gaussian blur JPEG compression Adversarial training
Near-OOD Far-OOD Near-OOD Far-OOD Near-OOD Far-OOD
SCALE 72.68 / 70.15 29.65 / 92.72 69.73 / 71.84 27.46 / 93.56 66.99 / 76.10 24.77 / 94.40
OptFS 71.84 / 69.66 30.27 / 92.24 69.43 / 71.12 28.34 / 92.95 70.11 / 73.51 28.32 / 93.42
AdaSCALE-A 67.53 / 75.18 28.11 / 93.31 61.24 / 77.59 25.87 / 94.06 60.21 / 80.68 21.13 / 95.35
[…]
Table 9: AdaSCALE-A with restricted access to ID data using ResNet-50 network in FPR@95 $\downarrow$ / AUROC $\uparrow$ format.
@@
ID statistics. With the rise of large models, where training data is often undisclosed or inaccessible, relying on full training ID datasets for OOD detection has become increasingly impractical. We rigorously assess AdaSCALE’s effectiveness with limited data by conducting experiments on ImageNet-1k with ResNet-50 using $n_{\text{val}}$ ID samples to compute ID statistics, where $n_{\text{val}}\in\{10,100,1000,5000\}$. The results in Table˜9 confirm that even with substantially restricted access to ID data, AdaSCALE-A achieves state-of-the-art performance.
@@
Latency. AdaSCALE incurs extra forward pass to compute perturbed activation $\textbf{a}^{\epsilon}$. Also, top-k operations (time complexity: $\mathcal{O}(D\log D)$) are applied to $Q$ and $C_{o}$ to estimate OODness. Comparing variable vs. fixed percentiles for scaling in Table˜10 over 10,000 trials, we observe that variable percentiles induce higher latency, though the latency ratio decreases with higher-dimensional activation spaces. While SCALE and OptFS have similar latency, AdaSCALE’s is 2.91x higher with gradient attribution for trivial pixel perturbation, dropping to 1.56x with random pixel perturbation.

### v3

@@ 4.2.1 Perturbation Mechanism
Let $x\in\mathbb{R}^{C_{\text{in}}\times H\times W}$ be an input image with $C_{\text{in}}$ input channels, $H$ height, and $W$ width. We denote channel value at position $(c,h,w)$ as $x[c,h,w]$. To identify channel values for perturbation, we employ pixel attribution that quantifies each input element’s influence on the model’s prediction. An attribution function, $AT(x,c,h,w)$, assigns a score to each channel value, with lower absolute scores indicating less influence. We select $o\%$ of channel value indices with lowest absolute attribution scores, forming the set $R$. We use a gradient-based attribution $AT(x,c,h,w)=\frac{\partial(g_{W}(f_{\theta}(x)))_{y_{pred}}}{\partial x[c,h,w]}$, where $y_{pred}$ is predicted class index. To create a perturbed input, we select a subset $R$ containing $o\%$ of channel values to perturb. If $\varepsilon$ is perturbation magnitude, perturbed image $x^{\varepsilon}$ is obtained as:
$x^{\varepsilon}[c,h,w]=\begin{cases}x[c,h,w]+\varepsilon\cdot\text{sign}(AT(x,c,h,w)),&\text{if }(c,h,w)\in R\\ x[c,h,w],&\text{if }(c,h,w)\notin R\end{cases}$
Remark: Is attribution necessary for perturbation? While we employ gradient-based attribution for principled pixel selection for perturbation, as we show later in Section˜F.5, it is important to note that even random selection empirically performs similarly, whereas selecting salient pixels degrades performance.
@@ 4.2.2 Activation Shift as OOD Indicator
Figure 3: $Q_{\text{OOD}}$$Q_{\text{ID}}$ vs $Q^{\prime}_{\text{OOD}}$$Q^{\prime}_{\text{ID}}$ in various OOD datasets with ResNet-50 on ImageNet-1k. $Q^{\prime}_{\text{OOD}}/Q^{\prime}_{\text{ID}}<Q_{\text{OOD}}/Q_{\text{ID}}$ suggests $C_{o}$ helps mitigate overconfident estimations.
Indeed, Figure˜3 illustrates that $Q^{\prime}_{\text{OOD}}/Q^{\prime}_{\text{ID}}<Q_{\text{OOD}}/Q_{\text{ID}}$, suggesting that the correction term $C_{o}$ helps mitigate overconfident estimations. If $\bar{Q}_{s}=\{\bar{Q}^{\prime}_{1},\bar{Q}^{\prime}_{2},...,\bar{Q}^{\prime}_{n_{\text{val}}}\}$ be the set of $Q^{\prime}$ values on $n_{\text{val}}$ ID validation samples, we could transform any $Q^{\prime}$ into a normalized probability scale by constructing empirical cumulative distribution function (eCDF) derived from $\bar{Q}_{s}$. The eCDF, denoted as $F_{Q^{\prime}}(Q^{\prime})$, can be defined as:
$F_{Q^{\prime}}(Q^{\prime})=\frac{1}{n_{\text{val}}}\sum_{i=1}^{n_{\text{val}}}\mathbbm{1}(\bar{Q}^{\prime}_{i}\leq Q^{\prime})$ (6)
where $\mathbbm{1}(\cdot)$ is the indicator function. A higher value of $F_{Q^{\prime}}(Q^{\prime})$ indicates a higher likelihood of the sample being OOD. Importantly, our experiments suggest that as few as 10 ID validation samples are sufficient to construct an effective eCDF for this purpose (See Table˜9).
@@ 4.3 Proposed Approach: Adaptive Scaling
Building on our observations, we propose AdaSCALE (Adaptive SCALE), a novel approach that introduces dynamic, sample-specific adjustments to the scaling procedure. The key insight is that $p^{th}$ percentile threshold should be a function of each test sample’s estimated OODness rather than a fixed value. The scaling factor $r$ increases as the $p^{th}$ percentile threshold rises (i.e., when more activations are excluded from the denominator in Equation 1). For optimal ID-OOD separation, we must scale ID samples more strongly than OOD samples, requiring a higher $p^{th}$ percentile for ID samples. We define an adaptive percentile threshold as:
$p=p_{\min}+(1-F_{Q^{\prime}}(Q^{\prime}))\cdot(p_{\max}-p_{\min})$ (7)
where $p_{\text{min}}$ and $p_{\text{max}}$ define the minimum and maximum limits of percentile threshold. It ensures samples with lower estimated OODness receive higher percentile thresholds, resulting in stronger scaling. (See Algorithm˜1). We implement two variants: AdaSCALE-A scales activations as $\mathbf{a}_{\text{scaled}}=\mathbf{a}\cdot\exp(r)$ (Djurisic et al., 2023; Xu et al., 2024). AdaSCALE-L scales logits as $\mathbf{z}_{\text{scaled}}=\mathbf{z}\cdot r^{2}$ (Djurisic et al., 2024). This approach is also outlined in Figure˜4.
@@ 5 Experiments
We use pre-trained models provided by PyTorch for ImageNet-1k experiments. For CIFAR experiments, we train three models per network using the standard cross-entropy loss and report the mean results across these three independent trials. The evaluation setup is provided in Table˜1. The non-percentile hyperparameters $(\lambda=10,k_{1}=1\%,k_{2}=5\%,o=5\%,\epsilon=0.5)$ were determined only once using ResNet-50 model via OpenOOD’s (Yang et al., 2022; Zhang et al., 2023c) automatic parameter search. We tune $(p_{\text{min}},p_{\text{max}})$ for optimal results in each case (see Appendix˜H). However, near-optimal performance can be achieved by tuning only $p_{\text{max}}$ while fixing $p_{\text{min}}$ to $60$ across all 22 cases (See Table˜5) – difference is trivial. Best results are bold, and second-best results are underlined.
@@
Baselines. MSP, EBO, ReAct, MLS, ASH, SCALE, BFAct, LTS, OptFS. See Section˜D.3 for additional comparisons. Currently, SCALE is the best-performing method (with ResNet-50), while OptFS is the best-generalizing method.
@@ 5.1 Empirical Results
[…]
CIFAR benchmark. We compare AdaSCALE with post-hoc baselines on CIFAR benchmarks using WRN-28-10 and DenseNet-101 networks, reporting the averaged performance in Table˜2. AdaSCALE outperforms all methods in average AUROC metric across CIFAR benchmarks in both near- and far-OOD detection. For far-OOD detection on CIFAR-10 benchmark, AdaSCALE-A achieves the best FPR@95 score of 33.11, outperforming the MSP baseline by approximately 1.4 points. Similarly, AdaSCALE-A attains the best FPR@95 / AUROC of 43.07 / 90.31 in near-OOD detection, though MSP remains competitive. In near-OOD detection on CIFAR-100 benchmark, AdaSCALE-A achieves the highest AUROC of 81.35, while in far-OOD detection, AdaSCALE-L reaches the best performance with FPR@95 / AUROC of 52.49 / 82.21. While activation-shaping methods perform well in ImageNet-1k, they seem to underperform in CIFAR. In contrast, AdaSCALE achieves consistently superior performance across all setups.
@@
ImageNet-1k benchmark. We compare our proposed method, AdaSCALE, with recent state-of-the-art approaches across eight architectures on the ImageNet-1k benchmark, as presented in Table˜3. AdaSCALE demonstrates consistently strong performance across all architectures compared to existing methods. Specifically, it surpasses the best-generalizing method, OptFS, by $\textbf{14.94\%}/\textbf{8.96\%}$ in the FPR@95/AUROC metric for near-OOD detection across all architectures. Additionally, it outperforms the best-performing method, SCALE (on ResNet-50), by $\textbf{12.96\%}/\textbf{6.44\%}$ in the same metric. A closer observation reveals that while OptFS excels in architectures such as EfficientNet, ViT-B-16, and Swin-B, scaling baselines perform comparably or even better in architectures like ResNet-50, ResNet-101, RegNet-Y-16, and DenseNet-201. In contrast, AdaSCALE-A achieves the best performance in near-OOD detection across all architectures, except for Swin-B, where BFAct performs optimally. Furthermore, effectiveness of AdaSCALE extends beyond near-OOD detection to far-OOD detection, demonstrating an average gain of 21.67% over OptFS in the FPR@95 metric.
Table 4: Avg. FSOOD detection on ImageNet-1k over 8 architectures.
Method Near-OOD Far-OOD ReAct 87.22 / 51.38 67.23 / 69.53 ASH 87.01 / 52.02 72.36 / 65.70 SCALE 86.75 / 52.27 69.36 / 68.97 BFAct 87.12 / 51.14 66.13 / 69.69 LTS 86.46 / 53.29 66.76 / 71.63 OptFS 85.83 / 52.17 63.32 / 71.44 AdaSCALE-A 81.34 / 55.03 58.87 / 72.41 AdaSCALE-L 81.62 / 55.14 59.19 / 72.85
FSOOD Detection. FSOOD detection extends conventional OOD detection by incorporating model’s ability to generalize on covariate-shifted ID inputs. We present FSOOD detection results in Table˜4 in FPR@95 $\downarrow$ / AUROC $\uparrow$ format. We can observe that this is a challenging task, as covariate-shifted ID datasets cause significant performance drop for all methods compared to conventional case. Yet, AdaSCALE outperforms OptFS by 4.49 and 4.13 points on average in FPR@95 for FSOOD detection across both near- and far-OOD datasets.
Accuracy: Like SCALE & LTS, AdaSCALE applies linear transformations to scale activations or logits, preserving accuracy, unlike post-hoc rectification methods (ASH, ReAct).
@@ 5.2 Ablation / Hyperparameter studies / Discussion
Constrained hyperparameter tuning. The hyperparameters $(\lambda,k_{1},k_{2},o,\epsilon)$ = $(10,1\%,5\%,5\%,0.5)$ determined with a ResNet-50 in ImageNet-1k benchmark generalize across architectures and datasets. To demonstrate this, we conduct experiments only allowing $p_{\text{max}}$ to be tuned across 8 and 2 architectures in ImageNet-1k and CIFAR-100 datasets, respectively and present the results in Table˜5. Inspired by early works that tune the percentile hyperparameter in $[60,99]$, we set $p_{\text{min}}$ to lower limit 60. Even under this highly constrained setting, AdaSCALE significantly outperforms the previous state-of-the-art (OptFS) by an average of 13% (FPR95) / 8% (AUROC) on near-OOD and 14% (FPR@95) / 2% (AUROC) on far-OOD datasets on ImageNet-1k benchmark. The superiority of AdaSCALE over OptFS is also evident in a small-scale setting (CIFAR-100).
Table 6: Robustness analysis with ResNet-50 in ImageNet-1k.
Metric / Method Gaussian blur JPEG compression Adversarial training
FPR@95 $\downarrow$ (Near-OOD / Far-OOD)
SCALE 72.68 / 29.65 69.73 / 27.46 66.99 / 24.77
OptFS 71.84 / 30.27 69.43 / 28.34 70.11 / 28.32
AdaSCALE-A 67.53 / 28.11 61.24 / 25.87 60.21 / 21.13
AUROC $\uparrow$ (Near-OOD / Far-OOD)
SCALE 70.15 / 92.72 71.84 / 93.56 76.10 / 94.40
OptFS 69.66 / 92.24 71.12 / 92.95 73.51 / 93.42
AdaSCALE-A 75.18 / 93.31 77.59 / 94.06 80.68 / 95.35
Robustness analysis. We evaluate the practical robustness of AdaSCALE under both real-world image corruptions and adversarial training. Specifically, we compare AdaSCALE against SCALE and OptFS on ImageNet-1k using a ResNet-50, applying Gaussian blur (kernel=5, $\sigma\in(0.01,0.5)$) and JPEG compression (50%) to the input images, without any hyperparameter re-tuning. In addition, we evaluate AdaSCALE on an adversarially trained ResNet-50 ($\epsilon=0.05$) using the same hyperparameters as in the standard setting. The results (Near-OOD / Far-OOD), shown in Table˜6, demonstrate that AdaSCALE consistently outperforms SCALE and OptFS, indicating robustness across both input perturbations and diverse training schemes.
@@
Sensitivity study. We study sensitivity of AdaSCALE-A hyperparameters with ResNet-50 on ImageNet-1k (Table˜7). The optimal perturbation magnitude is $\varepsilon=0.5$, close to RGB means and sufficient to disturb peak activations. Furthermore, setting $\lambda=10$ ensures activation shift $Q$ dominates the regularizer $C_{o}$. Optimal performance at $k_{1}=1\%$ confirms that shifts in peak activations are the most reliable OOD signal, while $k_{2}=5\%$ empirically performs best. $o=5\%$ yielding optimal result supports our hypothesis that perturbation should be trivial enough to only disturb peak activations. (See Appendices˜E and F for complete study.)
How effective is $Q$ without scaling? As shown in Figure˜2, $Q$ alone, without scaling, is not sufficiently discriminative. Its value lies in providing an independent and complementary OOD detection signal that is distinct from activation-pattern and logit-based information. For example, when $Q$ is directly used for near-OOD scoring on ImageNet-1k benchmark with ResNet-50, it achieves an FPR@95 $\downarrow$ / AUROC $\uparrow$ of $79.81/72.32$. In contrast, incorporating $Q$ into the scaling mechanism substantially improves performance to $59.43/78.14$. Hence, $Q$ alone is not effective.
Table 8: Comparison between $Q^{\prime}$ and $Q$ in average OOD detection performance on ImageNet-1k benchmark over 8 architectures.
Metric Near-OOD Far-OOD
$Q$ 68.90 / 78.07 42.63 / 89.18
$Q^{\prime}$ 61.17 / 80.50 30.11 / 92.73
Is $Q^{\prime}$ really needed instead of $Q$? Yes. AdaSCALE integrates three independent sources of OOD evidence – activation patterns, logit information, and activation shift – none of which serves as an oracle on its own. While $Q$ (activation shift) captures feature-level instability, relying on it alone without a balancing mechanism can lead to large scaling factors that overwhelm the informative OOD signal present in the logits. By incorporating the balancing term $C_{o}$, $Q^{\prime}$ appropriately regulates the influence of activation shift and preserves the contribution of logit information for OOD detection. This design choice is empirically validated on ImageNet-1k benchmark in Table˜8. (See Section˜D.1).
Why does AdaSCALE outperform OptFS? OptFS applies interval-specific activation scaling which yields modified logits used for OOD scoring. While this can improve separability on average, it implicitly assumes that activation reshaping is beneficial for every sample. For near-OOD inputs with ID-like activation patterns, this reshaping can distort informative original logit signals and lead to suboptimal OOD detection. (See Appendix˜J for dataset-specific results for OptFS on the SSB‑Hard datasets). AdaSCALE avoids this rigidity by adaptively fusing multiple, partially independent OOD cues allowing alternative signals to compensate when one becomes unreliable.
Table 9: Restricted access to ID data.
@@
ID statistics. With the rise of large models, where training data is often undisclosed or inaccessible, relying on full training ID datasets for OOD detection has become increasingly impractical. We rigorously assess AdaSCALE’s effectiveness with limited data by conducting experiments on ImageNet-1k with ResNet-50 using $n_{\text{val}}$ ID samples to compute ID statistics, where $n_{\text{val}}\in\{10,100,1000,5000\}$. The results (FPR@95 $\downarrow$ / AUROC $\uparrow$) in Table˜9 confirm that even with substantially restricted access to ID data, AdaSCALE-A achieves state-of-the-art performance.
@@
Latency. AdaSCALE requires an additional forward pass to compute the perturbed activation $\mathbf{a}^{\epsilon}$ only when gradient-based pixel perturbation is used; random perturbation, does not incur this cost. Also, top-k operations (time complexity: $\mathcal{O}(D\log D)$) are applied to $Q$ and $C_{o}$ to estimate OODness. Comparing variable vs. fixed percentiles for scaling in Table˜10 over 10,000 trials, we observe that variable percentiles induce higher latency, though the latency ratio decreases with higher-dimensional activation spaces. While SCALE and OptFS have similar latency, AdaSCALE’s is 2.91x higher with gradient attribution for trivial pixel perturbation, dropping to 1.56x with random pixel perturbation.